OpenStudy (anonymous):

y=IxI+4 vertex? axis of symmetry? slopes? which way it opens?? PLEASE HELP

OpenStudy (wolfe8):

Do you know how to sketch it? I think that's the easiest way to answer all that

OpenStudy (anonymous):

barely? can u help?

OpenStudy (wolfe8):

Sure. You already have the equation in the form of y=mx+b where m is the slope and b is the gradient. So you have the slope and the gradient, can you sketch it now? Or, since it's a straight line, you can always take points x=any negative numbers, 0, and any positive numbers and plot it :)

OpenStudy (anonymous):

ok now I know how to sketch it, PLEASE help with the answer portion part....????

OpenStudy (anonymous):

starting with vertex

OpenStudy (wolfe8):

Alright so how does your graph look like?

OpenStudy (anonymous):

1 line slants right crosses 1 on x-axis I think

OpenStudy (anonymous):

the axis across

OpenStudy (wolfe8):

What about for when x is negative? Can you draw it here using the option below?

OpenStudy (anonymous):

ok two lines, drawed

OpenStudy (wolfe8):

Can I see it? So I can check if you did it right and I can show you how to answer the rest?

OpenStudy (wolfe8):

I'm not sure if that's how it's supposed to be. Let's do it together. Let's use the range -8 to 8 with increments of 2 for x :) So we have  Now can you sketch the graph?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

graphed it, still don't know vertex, axis of symmetry, slopes, and what way it opens though??

OpenStudy (wolfe8):

How does it look like? From the graph, vertex is the point where the graph is 'sharp', axis of symmetry is the line where the graph can be reflected symmetrically, and the what shape did you get? In what direction is it open or, forms a containment?

OpenStudy (anonymous):

can I post a pic of my graph and u help me with the answer portion of question??

OpenStudy (wolfe8):

Yes. That's what I asked you to do :)

OpenStudy (wolfe8):

I'm sorry I'm trying to help 3 people at once. Alright so at what point do you see the graph 'sharp' or forming a vertex?

OpenStudy (anonymous):

0,4????

OpenStudy (wolfe8):

Not quite, look carefully. The point where the two seemingly separate lines meet. The V

OpenStudy (anonymous):

1,2 ???

OpenStudy (wolfe8):

Yes :) Now, if you draw a vertical line at going through that vertex, what do you see? Does the left side mirror the right side?

OpenStudy (anonymous):

yes, but what points would I put for the axis of symmetry?

OpenStudy (wolfe8):

So an axis is a line. You can have point of symmetry(I think) but here you have a line. Since it is a vertical line, it contains all the y values for a specific x value, right? So you just put what the x value is there in the form x=a where a is the x value

OpenStudy (anonymous):

doesn't make sense?? waz the axis of symmetry?? slopes???

OpenStudy (wolfe8):

Read what I said again. Axis of symmetry is a line, where it's left side mirrors the right side.

OpenStudy (anonymous):

axis of symmetry is -2??

OpenStudy (wolfe8):

No. I told you the line goes through the vertex.

OpenStudy (anonymous):

idk what it is then

OpenStudy (anonymous):

cuz, I think she only wants one number??? idk ur confusing and still don't know answers )':

OpenStudy (wolfe8):

The line will have the same x value as the vertex. x=the x-value for your vertex

OpenStudy (anonymous):

still don't ge it

OpenStudy (anonymous):

is the axis o symmetry 1

OpenStudy (wolfe8):

Yes. x=1

OpenStudy (anonymous):

yes!!! I got it right!!!! thanks.. what are the slopes?

OpenStudy (anonymous):

????????

OpenStudy (wolfe8):

The slope can be found using the formula difference in y/difference in x

OpenStudy (anonymous):

y2-y1 / x2-x1 ??????????

OpenStudy (wolfe8):

Yes. Now pick any 2 points along the same line and put the values into the formula.

OpenStudy (anonymous):

thanks... help with 3 more??? they're easier...

OpenStudy (wolfe8):

I'm kind of tired...Maybe someone else can help you with those :) But you should be able to solve them with the same concept

OpenStudy (anonymous):

pleaseeee???? it will only take a few minutes?? I partly understand them even please

OpenStudy (wolfe8):

Sure I'll try but I might not reply

OpenStudy (anonymous):

W/out graphing identify -vertex, -axis of symmetry -slopes

OpenStudy (anonymous):

8) y=2Ix+1I 9.) y=Ix-2I+3 10.) y=-1/2IxI-2

OpenStudy (wolfe8):

http://www.mathwebb.com/a2_notes_2-7.pdf Would this help?